OpenStudy (anonymous):

What is 3,920,000,000,000 in scientific notation?

OpenStudy (anonymous):

u want me 2 answer it

OpenStudy (anonymous):

????? yes or no

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

jk srry

OpenStudy (anonymous):

idk the answer srry ;(

OpenStudy (anonymous):

How many zeros can you count in there?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

With big numbers like that it's easy to lose track, there are 10 of them.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

You can shrink those 10 zeroes into a scientific notation of \[10^{10}\]

OpenStudy (anonymous):

Only thing you need to add is the numbers that aren't zero.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[392 \times 10^{10}\]

OpenStudy (anonymous):

Each time you move a decimal point left you raise the exponent by one as well.

OpenStudy (anonymous):

\[39.2 \times 10^{11}\] \[3.92 \times 10^{12}\]

OpenStudy (aum):

\[ 3,920,000,000,000 = 3.92 \times 10^{12} \]

OpenStudy (anonymous):

I agree